# GraphCheck: Breaking Long-Term Text Barriers with Extracted Knowledge Graph-Powered Fact-Checking

### v1

@@ Abstract
Large language models (LLMs) are widely used, but they often generate subtle factual errors, especially in long-form text. These errors are fatal in some specialized domains such as medicine. Existing fact-checking with grounding documents methods face two main challenges: (1) they struggle to understand complex multihop relations in long documents, often overlooking subtle factual errors; (2) most specialized methods rely on pairwise comparisons, requiring multiple model calls, leading to high resource and computational costs. To address these challenges, we propose GraphCheck, a fact-checking framework that uses extracted knowledge graphs to enhance text representation. Graph Neural Networks further process these graphs as a soft prompt, enabling LLMs to incorporate structured knowledge more effectively. Enhanced with graph-based reasoning, GraphCheck captures multihop reasoning chains which are often overlooked by existing methods, enabling precise and efficient fact-checking in a single inference call. Experimental results on seven benchmarks spanning both general and medical domains demonstrate a 6.1% overall improvement over baseline models. Notably, GraphCheck outperforms existing specialized fact-checkers and achieves comparable performance with state-of-the-art LLMs, such as DeepSeek-V3 and OpenAI-o1, with significantly fewer parameters.111Our anonymous code is available at https://anonymous.4open.science/r/GraphCheck-1D43.
@@
Yingjian Chen1††thanks: Equal contributions, Haoran Liu211footnotemark: 1 , Yinhong Liu3, Rui Yang4, Han Yuan4, Yanran Fu1, Pengyuan Zhou5, Qingyu Chen6, James Caverlee2, Irene Li1††thanks: Corresponding author, 1University of Tokyo, 2Texas A&M University, 3University of Cambridge, 4Duke-NUS Medical School, 5Aarhus University, 6Yale University,
@@ 1 Introduction
Large language models (LLMs) Hurst et al. (2024); Dubey et al. (2024), have demonstrated powerful generative capabilities in various domains Yang et al. (2023); Lee et al. (2024); Liu et al. (2023a); Yang et al. (2024c). However, due to limitations in training data and the lack of integration of domain-specific knowledge, LLMs often “hallucinate” factual errors or inaccurate information McKenna et al. (2023); Zhang et al. (2023); Gao et al. (2023). As LLMs prioritize linguistic fluency and contextual relevance in their generation processes, the generated content may appear convincing while lacking factuality Huang et al. (2023); Ramprasad et al. (2024); Yang et al. (2024b). This issue is particularly concerning in specialized domains like medicine, where factual errors can compromise patient safety, leading to misdiagnoses, inappropriate treatments, and, in severe cases, life-threatening consequences Ahsan et al. (2023); Yang et al. (2024b). Therefore, ensuring the reliability and factual accuracy of LLM outputs is essential Liu et al. (2024c).
@@
We consider the task of fact-checking claims against grounding documents, where the goal is to assess factual consistency based on provided textual evidence Tang et al. (2024). Given the high cost and time demands of manual verification, modern fact-checking methods have shifted to automated approaches using LLMs or natural language inference (NLI) models Fu et al. (2023); Kim et al. (2024). Standard LLM-based checking methods take a straightforward approach by directly feeding documents and claims into LLM for fact-checking judgment (Figure 1, left). However, when dealing with long-form documents, they often struggle to capture complex entity relations and overlook subtle inconsistencies given large volumes of information. Additionally, long prompts may exceed the LLM’s context window, causing potential loss of relevant details and limiting the model from effective fact-checking. To address this, specialized methods Zha et al. (2023); Min et al. (2023); Liu et al. (2024b) decompose long documents into smaller chunks and break claims into atomic facts, enabling fine-grained evaluation at the price of computational cost and efficiency (Figure 1, middle).
@@
To address the problem of long text fact-checking, we propose GraphCheck (Figure 1, right), a graph-enhanced framework using extracted knowledge graphs (KGs) to capture multi-hop logical relations between entities, enhancing both global coherence and fine-grained understanding in long texts. We employ Graph Neural Networks (GNNs) to encode these graph structures and integrate the graph embeddings into LLM inputs. The direct comparison between the extracted document and claim graphs enables fine-grained factual verification in a LLM inference. The GNNs are trained on our curated general-domain synthetic graph data based on MiniCheck Tang et al. (2024) training set, while the LLMs remain frozen. Empirically, we find that despite being trained on general-domain data, our model achieves improved performance not only on general-domain datasets but also on medical-domain datasets, demonstrating that its graph-enhanced reasoning ability generalizes across domains. We also provide this dataset as a benchmark for future research, allowing the training and evaluation of graph-based fact-checking.
@@
Empirical Findings. We demonstrate the effectiveness and efficiency of GraphCheck, achieving a 6.1% improvement over the base model in fact-checking across extensive general and medical benchmarks.
@@ 2 Related Work and Background
Methods in Detecting Hallucination. Recent fact-checking research Yuan and Vlachos (2023); Kim et al. (2023) use Retrieval-Augmented Generation (RAG) Fan et al. (2024); Yang et al. (2025) and external knowledge bases like DBpedia Lehmann et al. (2015) and Wikidata Vrandečić and Krötzsch (2014) to verify generated claims by retrieving structured or semi-structured data. Another line of research Manakul et al. (2023); Mündler et al. (2023) focuses on verifying factual consistency using LLMs with grounding documents. These approaches harness LLMs’ reasoning and language capabilities to fact-check claims against textual evidence. While effective for short texts, they often fail to capture fine-grained inconsistencies in longer documents, limiting their accuracy. Our work builds on this second setting, aiming to improve fact-checking performance on long texts by enhancing LLMs with structured graph-based reasoning.
@@
Fact-Checking on Long Texts. To address the challenge of capturing detailed errors in long texts, recent methods have shifted towards using fine-grained units for fact-checking. Methods like FactScore Min et al. (2023), MiniCheck Tang et al. (2024), and ACUEval Wan et al. (2024) focus on extracting atomic units from the generated text to enable fine-grained fact verification. However, these fine-grained fact-checking methods often require multiple calls to verify each unit or triple, especially for long texts, which greatly increases computational cost and time. In contrast, our approach uses KGs to model complex entity relationships in long texts, enabling fine-grained verification in a single call. This avoids repetitive calls and significantly improves efficiency.
@@
Graph-based Methods for Enhancing Factuality. Previous graph-based fact-checking methods have primarily focused on isolated triple evaluations or document-level encoding, often overlooking the global graph structure and topological information. GraphEval Liu et al. (2024b) extracts triples from claims and evaluates their factual consistency individually using a pretrained NLI model. However, it also relies on pairwise comparisons and does not incorporate the overall graph structure, limiting its ability to capture complex relationships. FactGraph Ribeiro et al. (2022) employs graph encoders to process documents and summary semantic graphs extracted via OpenIE. It then combines text and graph embeddings through an MLP for the final prediction. However, as a pre-LLM method, it lacks the powerful contextual reasoning ability of modern models. AMRFact Qiu et al. (2024) leverages AMR graphs to represent document structures and guide factual summarization generation, focusing on structured summarization rather than direct fact verification. Unlike previous methods, our approach integrates a trainable GNN with an LLM, combining long-form contextual understanding with structured knowledge from extracted KGs. By incorporating graph reasoning, our model captures complex entity relationships and logical structures, enabling fine-grained fact verification in a single comparison. This enhanced reasoning ability allows the model to generalize effectively to specialized domains.
@@ 3 GraphCheck
In this section, we introduce our GraphCheck framework, designed for efficient fact-checking. Intuitively, GraphCheck first extracts structural information from KGs to enrich the input text, then leverages a LLM for verification. GraphCheck contains of three main steps: (1) Given a source document and a generated claim , we extract knowledge triples from them and construct corresponding KGs. (2) A trainable GNN encodes the entire graph, generating comprehensive graph embeddings. (3) These embeddings, along with the document and claim texts, are fed into a verifier LLM, with frozen parameters, enabling single-call fine-grained fact-checking with the help of structured graph information, as shown in Figure 2.
@@ 3.1 Graph Construction
To construct the KGs, we extract triples in the form of from the text, where each entity and relation captures key semantic information. To achieve this, a LLM is employed to automatically identify and extract these triples. The detailed prompt used for triple extraction is provided in Appendix F. Building on the extracted triplets, we construct a directed graph . Here, is the set of node (entity) features, where each denotes the feature vector for node . is the set of edge (relation) features, where denotes the edge feature vector for an edge from node to node . The node features and edge features from textual attributes are encoded using Sentence-Transformers.222https://huggingface.co/sentence-transformers/all-roberta-large-v1 For a given generated claim and its source document we extract the corresponding graphs and .
@@ 3.2 GraphCheck Verification
Text Encoding. For a given generated claim and the source document , we concatenate them following the verifying template shown in Appendix F, and pass the rendered prompt into the verifier LLM to obtain the text embedding .
@@ 4.1 Datasets
Training Dataset. To train a GNN for extracting KG information, we use the {claim, document, label} pairs from MiniCheck dataset Tang et al. (2024) with 14K synthetic samples. We use Claude-3.5-Sonnet Anthropic to extract KG triples for claims and documents, constructing graphs for each pair. The final training dataset is structured as {, , , , label}. The dataset is split into training, validation, and test sets in a 6:2:2 ratio for model training and evaluation.
@@ 4.2 Baselines
Large-scale LLMs333We consider Large-scale LLMs as models with more than 300 B parameters. include GPT-4 Achiam et al. (2023), GPT-4o Hurst et al. (2024), OpenAI o1 Jaech et al. (2024), Claude 3.5-Sonnet Anthropic , and the largest open source model DeepSeek-V3 671B Liu et al. (2024a). For small-scale LLMs, we include Llama3 8B, Llama3.3 70B Dubey et al. (2024), Qwen2.5 7B, and Qwen2.5 72B Yang et al. (2024a). For specialized fact-checking methods, we include AlignScore Zha et al. (2023) and fine-grained fact-checkers like MiniCheck Tang et al. (2024) and ACUEval Wan et al. (2024). Additionally, we also consider graph-based methods, namely GraphEval Sansford et al. (2024) and GraphRAG Edge et al. (2024).
@@ 4.3 Evaluation Metric
Considering the data imbalance in some benchmarks, models biased towards a particular class in predictions may not reflect their true performance. To address this, we follow previous approached Liu et al. (2023b); Tang et al. (2023) and calculate balanced accuracy (BAcc). For more implementation details, please refer to Appendix D.
@@ 5.1 Main Results
Table 2 presents the BAcc of our proposed method, GraphCheck, compared to LLMs and specialized fact-checkers across general and medical domain benchmarks. The results show that our proposed GraphCheck achieves strong performance, reaching an overall BAcc of 71.1% across all benchmarks. Specifically, among large-scale LLMs, Claude 3.5-Sonnet achieves the best overall performance. Our method outperforms GPT-4 and GPT-4o and comes close to the most advanced large-scale models, including OpenAI o1, Claude 3.5-Sonnet, and the latest open-source model DeepSeek-V3 671B, while operating at a smaller scale and significantly lower cost. Interestingly, GPT-4o underperforms on the medical datasets COVID-Fact and PubHealth, which contain shorter texts, even scoring lower than GPT-4. For small-scale LLMs, our method achieves a 5.8% and 7.5% improvement over its similarly sized models, Llama3.3 70B and Qwen2.5 72B, respectively. For Specialized Fact-checking Methods, GraphCheck outperforms all methods, achieving 10.5%, 3%, and 6% improvements over ACUEval, MiniCheck, and GraphEval, respectively. Notably, compared to methods that require multiple calls, our method achieves superior performance with a single model call. Although GraphRAG is not typically used for fact-checking, its popularity motivated us to adapt it for this purpose. A detailed analysis of these adaptations is provided in the Appendix G.
@@ 5.2 Ablation Studies
Impact of Additional Graph Information. To evaluate the effectiveness of we proposed incorporating graph information, we compare (1) the base LLM models with (2) our proposed GraphCheck, which is based on these models. As shown in Figure 4, our approach has a significant improvement on both lightweight models (Llama3 8B444https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, Qwen2.5 7B555https://huggingface.co/Qwen/Qwen2.5-7B-Instruct) and larger models (Llama3.3 70B, Qwen2.5 72B). Specifically, as shown in Table 1, our method achieves significant improvement on relatively long-text datasets AggreFact-XSum, AggreFact-CNN, and Summeval. In contrast, for shorter-text datasets like COVID-Fact, the improvement is minimal, especially on PubHealth, our method even shows a slight performance drop. This is because the base models already handle fact-checking effectively on shorter, simpler texts, and introducing additional graph information may interfere with their inference.
@@
Impact of Training Data Sizes. To evaluate the impact of training data size on the fact-checking performance of GraphCheck-Llama3.3 70B, we conducted experiments across all benchmarks, as shown in Figure 5. The results demonstrate a general upward trend in model performance as the amount of training data increases. Specifically, significant improvements can be observed on the long-text datasets AggreFact-XSum, AggreFact-CNN, Summeval, and SCIFact. Among them, XSum exhibits the largest improvement, increasing from 60.1% to 72.9%, while CNN and Summeval also achieve approximately 10% improvements. In contrast, for the short-text datasets, our method shows only a slight improvement on COVID-Fact, while on PubHealth, performance gradually declines. These results further validate the conclusion drawn in Section 5.2
@@
From the above results, we can observe that as the training data size increases, the overall model performance shows a upward trend. Therefore, we believe that further increasing the data size could continue to enhance the performance of our proposed GraphCheck.
@@
Impact of Generated Knowledge Graph Quality. Due to the inherent randomness in extracting entity-relationship triples from text using LLMs, we conducted an experiment to assess how the quality of KGs generated from text impacts the model’s final fact-checking results, as illustrated in Figure 6. For shorter generated claims, the triples extracted by the four models show minimal differences, except for occasional missing details (e.g., "Saturday") by GPT-4o and DeepSeek-V3. These missing have minor effects on fact-checking results, as the models also relied on the original text for verification. In contrast, for longer document texts, there are significant differences in the quality of the triples generated by the models. Specifically, the triplets extracted by the Llama 8B model lacked crucial details, such as the time ("Saturday") and the reason for the exam cancellation. The loss of key information could turn the KG into interference during fact-checking, ultimately leading to incorrect results. On the other hand, while the language expressions of the triples extracted by GPT-4, Claude 3.5, and DeepSeek-V3 are different, they all capture the essential details and still ensure the fact-checker makes the right prediction.
@@
These findings indicate that for short texts, the quality of the extracted KG has minimal impact on fact-checking performance, as models still rely on the original text for verification. However, for long-form documents, the completeness of the KG is critical. If the KG lacks key fact-checking information, it misleads the model rather than assists in verification. This is because longer texts make it more difficult for the model to extract essential details directly, increase its dependence on the KG. In such cases, an incomplete or inaccurate KG introduces noise and ultimately compromises fact-checking accuracy. Conversely, if the missing information is irrelevant to the verification process, its absence does not affect the result.

### v2

@@ Abstract
Large language models (LLMs) are widely used, but they often generate subtle factual errors, especially in long-form text. These errors are fatal in some specialized domains such as medicine. Existing fact-checking with grounding documents methods face two main challenges: (1) they struggle to understand complex multihop relations in long documents, often overlooking subtle factual errors; (2) most specialized methods rely on pairwise comparisons, requiring multiple model calls, leading to high resource and computational costs. To address these challenges, we propose GraphCheck, a fact-checking framework that uses extracted knowledge graphs to enhance text representation. Graph Neural Networks further process these graphs as a soft prompt, enabling LLMs to incorporate structured knowledge more effectively. Enhanced with graph-based reasoning, GraphCheck captures multihop reasoning chains that are often overlooked by existing methods, enabling precise and efficient fact-checking in a single inference call. Experimental results on seven benchmarks spanning both general and medical domains demonstrate up to a 7.1% overall improvement over baseline models. Notably, GraphCheck outperforms existing specialized fact-checkers and achieves comparable performance with state-of-the-art LLMs, such as DeepSeek-V3 and OpenAI-o1, with significantly fewer parameters.111Our code is available at https://github.com/Yingjian-Chen/GraphCheck.
@@
Yingjian Chen1††thanks: Equal contributions, Haoran Liu211footnotemark: 1 , Yinhong Liu3, Jinxiang Xie1, Rui Yang4, Han Yuan4, Yanran Fu1, Peng Yuan Zhou5, Qingyu Chen6, James Caverlee2, Irene Li1††thanks: Corresponding author, 1University of Tokyo, 2Texas A&M University, 3University of Cambridge, 4Duke-NUS Medical School, 5Aarhus University, 6Yale University. irene.li@weblab.t.u-tokyo.ac.jp
@@ 1 Introduction
Large language models (LLMs) Hurst et al. (2024); Dubey et al. (2024), have demonstrated powerful generative capabilities in various domains Yang et al. (2023); Lee et al. (2024); Liu et al. (2023a); Yang et al. (2024c). However, due to limitations in training data and the lack of integration of domain-specific knowledge, LLMs often “hallucinate” factual errors or inaccurate information McKenna et al. (2023); Zhang et al. (2023); Gao et al. (2023). As LLMs prioritize linguistic fluency and contextual relevance in their generation processes, the generated content may appear convincing while lacking factuality Huang et al. (2023); Ramprasad et al. (2024); Yang et al. (2024b). This issue is particularly concerning in specialized domains like medicine, where factual errors can compromise patient safety, leading to misdiagnoses, inappropriate treatments, and, in severe cases, life-threatening consequences Ahsan et al. (2023); Yang et al. (2024b). Therefore, ensuring the reliability and factual accuracy of LLM outputs is essential (Liu et al., 2024c).
@@
We consider the task of fact-checking claims against grounding documents, where the goal is to assess factual consistency based on provided textual evidence Tang et al. (2024b). Given the high cost and time demands of manual verification, modern fact-checking methods have shifted to automated approaches using LLMs or natural language inference (NLI) models Fu et al. (2023); Kim et al. (2024). Standard LLM-based checking methods take a straightforward approach by directly feeding documents and claims into LLM for fact-checking judgment (Figure 1, left). However, when dealing with long-form documents, they often struggle to capture complex entity relations and overlook subtle inconsistencies given large volumes of information. Additionally, long prompts may exceed the LLM’s context window, causing potential loss of relevant details and limiting the model from effective fact-checking. To address this, specialized methods Zha et al. (2023); Min et al. (2023); Liu et al. (2024b) decompose long documents into smaller chunks and break claims into atomic facts, enabling fine-grained evaluation at the price of computational cost and efficiency (Figure 1, middle).
@@
To address the problem of long text fact-checking, we propose GraphCheck (Figure 1, right), a graph-enhanced framework using extracted knowledge graphs (KGs) to capture multi-hop logical relations between entities, enhancing both global coherence and fine-grained understanding in long texts. We employ Graph Neural Networks (GNNs) Veličković et al. (2017); Yun et al. (2019) to encode these graph structures and integrate the graph embeddings into LLM inputs He et al. (2024); Tian et al. (2024); Jin et al. (2024). The direct comparison between the extracted document and claim graphs enables fine-grained factual verification in an LLM inference. The GNNs are trained on our curated general-domain synthetic graph data based on MiniCheck Tang et al. (2024b) training set, while the LLMs remain frozen. Empirically, we find that despite being trained on general-domain data, our model achieves improved performance not only on general-domain datasets but also on medical-domain datasets, demonstrating that its graph-enhanced reasoning ability generalizes across domains. We also provide this dataset as a benchmark for future research, allowing the training and evaluation of graph-based fact-checking.
@@
Empirical Findings. We demonstrate the effectiveness and efficiency of GraphCheck, achieving a 7.1% improvement over the base model in fact-checking across extensive general and medical benchmarks.
@@ 2 Related Work and Background
Methods in Detecting Hallucination. Recent fact-checking research Yuan and Vlachos (2023); Kim et al. (2023) uses Retrieval-Augmented Generation (RAG) Fan et al. (2024); Yang et al. (2025) and external knowledge bases like DBpedia Lehmann et al. (2015) and Wikidata Vrandečić and Krötzsch (2014) to verify generated claims by retrieving structured or semi-structured data.
@@
Another line of research Manakul et al. (2023); Mündler et al. (2023) focuses on verifying factual consistency using LLMs with grounding documents. These approaches harness LLMs’ reasoning and language capabilities to fact-check claims against textual evidence. While effective for short texts, they often fail to capture fine-grained inconsistencies in longer documents, limiting their accuracy. Our work builds on this second setting, aiming to improve fact-checking performance on long texts by enhancing LLMs with structured graph-based reasoning.
@@
Fact-Checking on Long Texts. To address the challenge of capturing detailed errors in long texts, recent methods have shifted towards using fine-grained units for fact-checking. Methods like FactScore Min et al. (2023), MiniCheck Tang et al. (2024b), and ACUEval Wan et al. (2024) focus on extracting atomic units from the generated text to enable fine-grained fact verification. However, these fine-grained fact-checking methods often require multiple calls to verify each unit or triple, especially for long texts, which greatly increases computational cost and time. In contrast, our approach uses KGs to model complex entity relationships in long texts, enabling fine-grained verification in a single call. This avoids repetitive calls and significantly improves efficiency.
@@
Graph-based Methods for Enhancing Factuality. Previous graph-based fact-checking methods have primarily focused on isolated triple evaluations or document-level encoding, often overlooking the global graph structure and topological information. GraphEval Liu et al. (2024b) extracts triples from claims and evaluates their factual consistency individually using a pretrained NLI model. However, it also relies on pairwise comparisons and does not incorporate the overall graph structure, limiting its ability to capture complex relationships. FactGraph Ribeiro et al. (2022) employs graph encoders to process documents and summary semantic graphs extracted via OpenIE. It then combines text and graph embeddings through an MLP for the final prediction. However, as a pre-LLM method, it lacks the powerful contextual reasoning ability of modern models. AMRFact Qiu et al. (2024) leverages AMR graphs to represent document structures and guide factual summarization generation, focusing on structured summarization rather than direct fact verification. Unlike previous methods, our approach integrates a trainable GNN with an LLM Tian et al. (2024); He et al. (2024), combining long-form contextual understanding with structured knowledge from extracted KGs. By incorporating graph reasoning, our model captures complex entity relationships and logical structures, enabling fine-grained fact verification in a single comparison. This enhanced reasoning ability allows the model to generalize effectively to specialized domains. An extended discussion on how LLMs understand knowledge graphs is provided in Appendix F.
@@ 3 GraphCheck
In this section, we introduce our GraphCheck framework, which is designed for efficient fact-checking. Intuitively, GraphCheck first extracts structural information from KGs to enrich the input text and then leverages an LLM for verification. GraphCheck contains three main steps: (1) Given a source document and a generated claim , we extract knowledge triples from them and construct corresponding KGs. (2) A trainable GNN encodes the entire graph, generating comprehensive graph embeddings. (3) These embeddings, along with the document and claim texts, are fed into a verifier LLM, with frozen parameters, enabling single-call fine-grained fact-checking with the help of structured graph information, as shown in Figure 2.
@@ 3.1 Graph Construction
To construct the KGs, we extract triples in the form of from the text, where each entity and relation captures key semantic information. To achieve this, an LLM is employed to automatically identify and extract these triples. The detailed prompt used for triple extraction is provided in Appendix H. Building on the extracted triplets, we construct a directed graph . Here, is the set of node (entity) features, where each denotes the feature vector for node . is the set of edge (relation) features, where denotes the edge feature vector for an edge from node to node . The node features and edge features from textual attributes are encoded using Sentence-Transformers.222https://huggingface.co/sentence-transformers/all-roberta-large-v1 For a given generated claim and its source document we extract the corresponding graphs and .
@@ 3.2 GraphCheck Verification
Text Encoding. For a given generated claim and the source document , we concatenate them following the verifying template shown in Appendix H, and pass the rendered prompt into the verifier LLM to obtain the text embedding .
@@ 4.1 Datasets
Training Dataset. To train a GNN for extracting KG information, we use the {claim, document, label} pairs from MiniCheck dataset Tang et al. (2024b) with 14K synthetic samples. We use Claude-3.5-Sonnet Anthropic to extract KG triples for claims and documents, constructing graphs for each pair. The final training dataset is structured as {, , , , label}. The dataset is split into training, validation, and test sets in a 6:2:2 ratio for model training and evaluation.
@@ 4.2 Baselines
Large-scale LLMs333We consider Large-scale LLMs as models with more than 300 B parameters. include GPT-4 OpenAI (2023), GPT-4o Hurst et al. (2024), OpenAI o1 Jaech et al. (2024), Claude 3.5-Sonnet Anthropic , and the largest open source model DeepSeek-V3 671B Liu et al. (2024a). For small-scale LLMs, we include Llama3 8B, Llama3.3 70B Dubey et al. (2024), Qwen2.5 7B, and Qwen2.5 72B Yang et al. (2024a). For specialized fact-checking methods, we include AlignScore Zha et al. (2023) and fine-grained fact-checkers like MiniCheck Tang et al. (2024b) and ACUEval Wan et al. (2024). Additionally, we also consider graph-based methods, namely GraphEval Sansford et al. (2024) and GraphRAG Edge et al. (2024).
@@ 4.3 Evaluation Metric
Considering the data imbalance in some benchmarks, models biased towards a particular class in predictions may not reflect their true performance. To address this, we follow previous approached Liu et al. (2023b); Tang et al. (2023) and calculate balanced accuracy (BAcc). For more implementation details, please refer to Appendix G.
@@ 5.1 Main Results
Table 2 presents the BAcc of our proposed method, GraphCheck, compared to LLMs and specialized fact-checkers across general and medical domain benchmarks. The results show that our proposed GraphCheck achieves strong performance, reaching an overall BAcc of 71.1% across all benchmarks. Specifically, among large-scale LLMs, Claude 3.5-Sonnet achieves the best overall performance. Our method outperforms GPT-4 and GPT-4o and comes close to the most advanced large-scale models, including OpenAI o1, Claude 3.5-Sonnet, and the latest open-source model DeepSeek-V3 671B, while operating at a smaller scale and significantly lower cost. Interestingly, GPT-4o underperforms on the medical datasets COVID-Fact and PubHealth, which contain shorter texts, even scoring lower than GPT-4. For small-scale LLMs, our method achieves improvements of 5.8% and 7.1% over the similarly sized base models, Llama3.3 70B and Qwen2.5 72B, respectively. For Specialized Fact-checking Methods, GraphCheck outperforms all methods, achieving 10.5%, 3%, and 6% improvements over ACUEval, MiniCheck, and GraphEval, respectively. Notably, compared to methods that require multiple calls, our method achieves superior performance with a single model call. Although GraphRAG is not typically used for fact-checking, its popularity motivated us to adapt it for this purpose. A detailed analysis of these adaptations is provided in the Appendix E.
@@
| Llama3.3 70B | 60.1 | 53.5 | 57.6 | 54.3 | 69.0 | 85.7 | 76.9 | 65.3 | |
| Llama3.3 70B + KG Text | 60.6 | 53.6 | 57.8 | 54.4 | 71.2 | 88.8 | 78.5 | 66.4 | |
| GraphCheck-Llama3.3 70B + KG Text | 66.3 | 60.2 | 63.8 | 55.6 | 71.2 | 87.9 | 70.4 | 67.9 | |
| GraphCheck-Llama3.3 70B | 72.9 | 62.4 | 67.3 | 60.3 | 71.5 | 89.4 | 73.6 | 71.1 |
@@ 5.2 Ablation Studies
Impact of Additional Graph Information. To evaluate the effectiveness of incorporating graph information, we compare (1) the base LLM models with (2) our proposed GraphCheck, which is based on these models. As shown in Figure 4, our approach has a significant improvement on both lightweight models (Llama3 8B444https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, Qwen2.5 7B555https://huggingface.co/Qwen/Qwen2.5-7B-Instruct) and larger models (Llama3.3 70B, Qwen2.5 72B). Specifically, as shown in Table 1, our method achieves significant improvement on relatively long-text datasets AggreFact-XSum, AggreFact-CNN, and Summeval.
| Method | COVID -Fact | Pub Health | MiniCheck TestSet | Reveal | Overall Avg. (%) |
| Qwen2.5 72B | 69.9 | 76.7 | 69.9 | 85.9 | 75.6 |
| Llama3.3 70B | 69.0 | 76.9 | 71.0 | 86.8 | 75.9 |
| AlignScore | 66.5 | 64.4 | 71.9 | 85.3 | 72.0 |
| MiniCheck | 65.9 | 66.3 | 88.8 | ||
| GraphEval | 60.7 | 63.7 | 72.1 | 89.8 | 71.6 |
| GraphCheck- Qwen2.5 72B | 69.7 | 73.6 | 78.4 | 88.0 | 77.4 |
| GraphCheck- Llama3.3 70B | 71.5 | 71.7 | 81.2 | 89.7 | 78.5 |
@@
Evaluating the Use of KGs as Prompt Text. Given that extracted KGs can be directly used as the prompt text for LLMs, as employed in G-Retriever He et al. (2024), we conduct ablation studies to assess the impact of using the knowledge graph as the prompt text for LLMs, as shown in Table 3. Experimental results show that directly adding KGs text into the prompt yields only marginal improvement over the base model Llama3 70B and even results in performance degradation when integrated into our proposed GraphCheck. This is mainly because LLMs struggle to capture fine-grained factual errors directly from the long textual inputs, and the large extracted knowledge graph text may further exacerbate this issue. In contrast, our proposed GraphCheck only integrates KGs through a GNN, which enables more effective integration of structured graph information.
Evaluation on Short-Text Fact-Checking. Although GraphCheck is specifically designed for long-term text fact-checking, we also provide evaluations on short texts (length < 150). To do so, we conducted additional experiments on four short-text datasets, as shown in Table 4. The MiniCheck TestSet refers to the test split obtained from the 14K dataset using a 6:2:2 ratio (see Section 4.1). The results demonstrate that our proposed GraphCheck maintains robust performance on short-text fact-checking tasks, outperforming existing specialized methods and base models, indicating its effectiveness across varying text lengths. Although the performance of GraphCheck shows degradation only on the PubHealth dataset compared to the base models, this is primarily because Qwen2.5 72B and Llama-3.3 70B already achieve exceptionally high performance on this dataset, even surpassing OpenAI’s o1 (as shown in Table 2).
@@
Impact of Training Data Sizes. To evaluate the impact of training data size on the fact-checking performance of GraphCheck-Llama3.3 70B, we conducted experiments across all benchmarks, as shown in Figure 5. The results demonstrate a general upward trend in model performance as the amount of training data increases. Specifically, significant improvements can be observed on the long-text datasets AggreFact-XSum, AggreFact-CNN, Summeval, and SCIFact. Among them, XSum exhibits the largest improvement, increasing from 60.1% to 72.9%, while CNN and Summeval also achieve approximately 10% improvements. In contrast, for the short-text datasets, our method shows only a slight improvement on COVID-Fact, while on PubHealth, performance gradually declines. These results further validate the conclusion drawn in Section 5.2.
@@
From the above results, we can observe that as the training data size increases, the overall model performance shows an upward trend. Therefore, we believe that further increasing the data size could continue to enhance the performance of our proposed GraphCheck. Additional experiments on different graph-building methods and GNN architectures are provided in Appendix D.
@@
Impact of Generated Knowledge Graph Quality. Due to the inherent randomness in extracting entity-relationship triples from text using LLMs, we conducted an experiment to assess how the quality of KGs generated from text impacts the model’s final fact-checking results, as illustrated in Figure 6. For shorter generated claims, the triples extracted by the four models show minimal differences, except for occasional missing details by GPT-4o and DeepSeek-V3. These missing have minor effects on fact-checking results, as the models also relied on the original text for verification. In contrast, for longer document texts, there are significant differences in the quality of the triples generated by the models. Specifically, the triplets extracted by the Llama 8B model lacked crucial details, such as the time ("Saturday") and the reason for the exam cancellation. The loss of key information could potentially turn the KG into interference during fact-checking, ultimately leading to incorrect results. On the other hand, while the language expressions of the triples extracted by GPT-4, Claude 3.5, and DeepSeek-V3 are different, they all capture the essential details and still ensure that the fact-checker makes the right prediction.
@@
These findings indicate that for short texts, the quality of the extracted KG has minimal impact on fact-checking performance, as models still rely on the original text for verification. However, for long-form documents, the completeness of the KG is critical. If the KG lacks key fact-checking information, it misleads the model rather than assists in verification. This is because longer texts make it more difficult for the model to extract essential details directly, increasing its dependence on the KG. In such cases, an incomplete or inaccurate KG introduces noise and ultimately compromises fact-checking accuracy. Conversely, if the missing information is irrelevant to the verification process, its absence does not affect the result.